Tickets to the school’s dinner theatre cost $6 for children and $9 for adults. 225 tickets were sold for a total of $1875
how many children attended?
how many adults attended

Answer by Maths68(1474)

Tickets to the school’s dinner theatre cost $6 for children and $9 for adults. 225 tickets were sold for a total of $1875
how many children attended?
how many adults attended
Let
Children attended = x
Adults attended = 225-x
Cost of Children's ticket = 6
Cost of Adult's ticket = 9
Total Amount of all tickets sold = 1875

(Children)(Cost of Children's ticket)+(Adults)(Cost of Adult's ticket)= Total Amount of all tickets sold
(x)(6)+(225-x)(9)=1875
6x+2025-9x=1875
-3x=1875-2025
-3x=-150
Divide by -3 both sides
-3x/-3=-150/-3
x=50

Children attended = x = 50
Adults attended = 225-x = 225-50=175